40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Feb 4 01:03:47 2011, Frank, DailyProgress, BEAT, cable delay setup cable-delay.png
    Reply  Fri Feb 4 13:59:48 2011, Frank, DailyProgress, BEAT, trend of the setup Untitled.png
Message ID: 471     Entry time: Fri Feb 4 13:59:48 2011     In reply to: 470
Author: Frank 
Type: DailyProgress 
Category: BEAT 
Subject: trend of the setup 

 This setup has too much gain (i.e. not enough range). Please reduce the arm length asymmetry by a factor of 10 so that we can monitor over 24 hours.

Also the temperature channels ought to be calibrated (via the EPICS .db) so that the readout is in degC instead of ARB.

Attachment 1: Untitled.png  52 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-